Circular No. 5604

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET TUTTLE (1992r) G. Tancredi and M. Lindgren report their recovery of this comet on

CCD exposures with the Nordic Optical Telescope, La Palma. The object is of stellar appearance, and the magnitudes are V.

1992 UT  α_{2000}  δ_{2000}  m_{2}
July 29.06250  20^{h}05^{m}39.^{s}19  +31^{o}35′56′′.8  21.2
29.14236  20 05 36.12  +31 35 59.6  21.3
29.97188  20 05 01.70  +31 35 49.9  21.4
30.89896  20 04 23.08  +31 35 30.9  21.2
31.19896  20 04 10.51  +31 35 23.5  21.4

The prediction on MPC 18257 requires correction by ΔT = −1.7 days. The following orbital elements, by the undersigned, satisfy 49 observations 1980-1992 with mean residual 1′′.0:

Epoch = 1994 June 17.0 TT

T = 1994 June 25.2907 TT  ω = 206.^{o}7030
e = 0.824089  Ω = 270.5485 }2000.0
q = 0.997732 AU  i = 54.6923

a = 5.671812 AU n^{o} = 0.0729661 P = 13.508 years

1992 TT α_{2000} δ_{2000} Δ r ϵ β m_{2}
Aug. 26 19^{h}47.^{m}58 +30^{o}37.^{′}5 5.504 6.140 125.^{o}0 7.^{o}7 21.2
Sept. 5 19 42.49 +29 53.8 5.507 6.085 120.8 8.2 21.2
15 19 38.53 +29 01.9 5.527 6.029 115.5 8.7 21.2
25 19 35.82 +28 04.9 5.561 5.973 109.6 9.1 21.2
Oct. 5 19 34.45 +27 05.7 5.606 5.916 103.3 9.5 21.2
15 19 34.42 +26 07.0 5.659 5.859  96.7 9.7 21.2
25 19 35.67 +25 11.4 5.716 5.801  90.0 9.9 21.2
Nov. 4 19 38.15 +24 20.9 5.774 5.743  83.3 9.9 21.2

FG SAGITTAE J. Papousek, Brno University Observatory, reports: “In contrast to

the relatively small light changes over V = 9.0-9.3 during the past 20 years, recent photometry shows this variable to be sharply declining in brightness: Sept. 2.88 UT, V = 10.05±0.01, B −V = +1.80±0.01.” Visual magnitude estimates by B. H. Granslo, Fjellhamar, Norway (also provided by the AAVSO): Aug. 7.03 UT, 9.2; 19.00, 9.3; 24.91, 9.4; 31.95, 9.6.

1992 September 4 Brian G. Marsden